# Constraint Breeds Generalization: Temporal Dynamics as an Inductive Bias

### v1

@@ 2.4 Experiment 3: Zero-Shot Transfer to Unseen Physical Regimes
The preceding experiments established that dissipative constraints embedded in input encoding (modulated by the parameter $\delta$) induce generalization at both representational (Exp 1) and structural (Exp 2) levels. A natural question arises: is this principle encoding-specific, or can equivalent constraints be realized through architectural parameters? To address this, we designed a two-fold validation targeting “unseen physical regimes” in reinforcement learning (RL) tasks[packer2018assessing, kirk2023survey]. Unlike standard approaches relying on extrinsic domain randomization[tobin2017domain, peng2018sim], we posit that optimal dynamical constraints confer intrinsic robustness regardless of implementation level.
  • Encoding-Level. We first verified that encoding-level constraints translate to behavioral robustness. Using CartPole [towers2024gymnasium] within a zero-shot paradigm, agents were trained on a standard “Easy” environment and evaluated on progressively difficult physical parameters (Appendix A.3.1). Dynamical encoded (duffing) states were processed by either SNNs or MLPs to isolate the role of temporal integration.
@@
  • Architecture-Level. While dynamical encoding serves as an effective probe for investigating how input constraints influence learning, it relies on synthetically expanding state vectors into trajectories. However, RL environments provide inputs that are inherently sequential. To demonstrate that the constraint principle generalizes beyond artificial injection, we directly varied the SNN’s membrane leak parameter $\beta$, which governs intrinsic temporal integration. This parallels our encoding-level analysis: just as $\delta$ controls phase space contraction in input trajectories, $\beta$ controls information dissipation within the network’s hidden states. We conducted systematic $\beta$-sweeps across CartPole (4D state, REINFORCE) and LunarLander [towers2024gymnasium] (8D state, PPO [schulman2017proximal]), comparing Leaky SNNs against MLP and LSTM baselines on raw state histories without external encoding (see Appendix A.3.2 for detailed implementation).
@@
Both validation approaches reveal consistent patterns (Tables 23): At the encoding level, SNN-Transition ($\delta=2.0$) achieved the lowest generalization gap (38.6 under fixed-budget; 42.4 with sufficient-training), significantly outperforming Expansive (87.6), Dissipative (53.1), and MLP Baseline (112.9). At the architecture level, a consistent optimum emerges: on CartPole, intermediate dissipation ($\beta=0.5$) achieved the lowest gap (13.4), a 90% improvement over MLP (138.7); on the more complex LunarLander, the optimal regime persisted at $\beta=0.5$ (Gap: 96.5), outperforming MLP (180.3) by 46% and LSTM (193.1) by 50%. We also observed a non-monotonic performance profile in SNNs across the parameter spectrum: over-constrained regimes ($\beta\leq 0.3$) exhibited training instability, while unconstrained SNNs ($\beta=1.0$) showed degraded generalization in complex tasks despite successful training. This cross-implementation convergence underscores that the efficacy of dynamical constraints is realization-agnostic: the optimal regime emerges from a characteristic timescale that optimally balances signal preservation against noise suppression.
@@
Table 2: Encoding-level generalization comparison (CartPole). Dynamical encoding transforms states into temporal trajectories with varying constraint strength ($\delta$). SNN-Transition ($\delta=2.0$) achieves the lowest gap, demonstrating that SNN temporal integration is required to decode constraint-induced invariants. Bold: best generalization (lowest gap).
@@
Table 3: Architecture-level generalization comparison. Varying membrane leak $\beta$ on raw state inputs (no external encoding) reveals a non-monotonic optimum: intermediate dissipation ($\beta\approx 0.5$) yields best generalization. Note that over-constrained regimes ($\beta\leq 0.3$) fail to converge reliably. Bold: best generalization (lowest gap).
@@
Furthermore, the non-monotonic performance profile observed in Table 3 helps disentangle a critical confound: does robustness originate from the temporal constraint itself, or merely from the “memory” capacity inherent in SNN membrane dynamics[eshraghian2023training]? If memory alone drove generalization, performance should monotonically increase with information retention (high $\beta$). Instead, the observed degradation (or failure) at $\beta\approx 1.0$ indicates that intermediate dissipation implements a necessary temporal information bottleneck, filtering transient noise while preserving invariant structure. This interpretation is further reinforced by evidence from recurrent variants (RLeaky SNNs, see Appendix 16): despite possessing additional memory pathways, these networks did not exhibit improved generalization. On the contrary, they displayed a significantly narrower stability window, failing to converge in unconstrained regimes ($\beta\approx 1.0$) where feedforward Leaky SNNs still functioned. This fragility suggests that excess memory without sufficient constraint destabilizes learning, revealing an inherent duality of the dynamical constraint, as it simultaneously governs both information retention (memory) and phase space contraction (forgetting).
In conclusion, these multi-level validations provide converging evidence for our central thesis: temporal constraint breeds generalization. Structured dissipation functions not as a limitation but as a temporal inductive bias, compelling networks to extract robust, invariant representations. This mechanism proves implementation-agnostic, whether constraints are imposed externally via (specific) encoding or internally via architecture.
@@ 3 Discussion
Related Work Recent literature increasingly frames physical constraints not as computational limitations, but as essential inductive biases for intelligence. These constraints can be implemented at different system levels: through architectural design[hernandez2022thermodynamics], learnable internal dynamics[zhang2024intrinsic], or temporal hierarchies[cheng2020finite, moro2024role]. Our work advances this frontier by identifying a foundational insight: temporal dissipation functions as an inductive bias, effective whether imposed via external input dynamics or internal architectural specification. Mechanistically, this process suppresses high-frequency fluctuations to favor stable trajectories, conceptually aligning with Slow Feature Analysis (SFA)[wiskott2002slow], which posits that invariance emerges from slowly varying signals. However, unlike SFA which requires an explicit optimization objective, we demonstrate that dissipative dynamics implicitly impose spectral compression as a physical realization of temporal invariance. Together, our results indicate that temporal constraints alone can induce the invariant representations required for out-of-distribution generalization[deng2022strong, goyal2022inductive].
@@
From Generalization to Specialization This framework provides a theoretical justification for the "performance valley" discussed earlier[chen2025dynamical]. To rigorously ground this observation, we formalize this trade-off phenomenon through a complementary PAC-Bayes analysis[mcallester1998some] (Appendix A.4): By bounding the generalization error as a function of the posterior’s alignment with the prior, this analysis reveals that the transition regime minimizes the trade-off between empirical fit (training error) and structural stability (Kullback–Leibler divergence) while exhibiting the most stable gradient norms (Appendix Table 1819). This theoretical insight re-interprets the observed performance dip as a necessary cost of abstraction, i.e., the increased training effort: as evidenced by the delayed convergence in our behavioral tasks (Exp 3, also in Appendix Figure 7), extracting robust invariants is inherently more computationally demanding than the rapid memorization of transient patterns. In other words, the very regime that appears suboptimal from a purely efficiency-centric view is proven to be globally optimal for learning stable, transferable solutions.
@@
This finding not only aligns with the "Edge of Chaos" theory of computational capacity [langton1990computation, bertschinger2004edge] but also extends its theoretical scope by explicitly identifying state temporal constraints as an active regularization mechanism. Taken together, we envision a different landscape: rather than seeking a static optimum for all computation, the neural computation landscape should be viewed as a continuous dynamical spectrum, where the transition regime acts as an adaptability reservoir, maintaining the scale-invariant complexity required to extract robust invariants. Moving away from this center, adjacent regimes serve distinct downstream objectives: shifting towards expansive dynamics ($\Sigma\lambda_{i}>0$) maximizes discriminability for inference, or collapsing towards dissipative dynamics ($\Sigma\lambda_{i}\ll 0$) ensures stability and energy minimization.
@@
Therefore, this implies that the neural system should be formalized not as a static entity, but as an adaptive dynamical system, where the “optimal” configuration is not a fixed point but a continuous spectrum through phase space. From this perspective, the neuron’s dynamical properties are not intrinsic constants but mutable state variables that evolve in response to environmental demands and internal budgets. For machine learning, this necessitates a shift from optimal architectural search to system dynamic regime navigation: robust intelligence emerges from "mortal" systems that possess the agency to modulate their internal states over time. In this context, temporal factor acts not only as a computational resource, but also as the fundamental balance between retaining specific data fidelity and extracting generalizable understanding. The transition regime plays the role of a generalization optimum, which emerges not from maximizing memory, but from aligning the integration timescale to task-relevant invariants. This perspective offers a potential explanation to the apparent paradox that biological neural systems, despite vast memory capacity or energy efficiency, exhibit robust generalization: sparse coding and metabolic constraints may represent evolutionary solutions that naturally situate neural dynamics within the transition regime.
@@
Limitations and Future Directions While this work demonstrates the efficacy of dynamical constraints, our implementation primarily employs a fixed oscillator system as a probe for mechanistic discovery: By isolating specific dynamical regimes, we establish a controlled environment for qualitative analysis, yet this represents only a static instantiation; the performance reported herein likely constitutes a lower bound on the paradigm’s potential. Although our architecture-level experiments ($\beta$-sweeps) demonstrate that the constraint principle generalizes beyond dynamical encoding, both implementations employ only fixed parameters. Meanwhile, biological systems exhibit heterogeneous time constants across neural populations; exploring learned or adaptive $\beta$ distributions represents a natural extension. The critical observation is that the transition regime is defined not by a specific parameter value, but by a spectral signature: the combination of high entropy (preserved information complexity) and low dominant frequency (alignment with neural networks’ spectral bias). As demonstrated in our scale-space analysis (Fig.5B), this signature remains stable across timescales ($T_{\max}$) and sampling densities ($N$), suggesting an intrinsic property of the dynamical regime rather than a parameter-tuning artifact. Nevertheless, we acknowledge that verifying this regime-generality across alternative dynamical systems remains an important direction for future work. Ultimately, this study suggests that the path toward robust AI lies not just in scaling model parameters, but in mastering the temporal constraints that shape the learning landscape itself.
@@ 4 Conclusion
This work identifies dynamical constraints as a distinct inductive bias operating on the temporal dimension. By optimally filtering transient noise while preserving signal structure, this regime induces robust invariance. We show that this capability is architecturally co-dependent: it requires temporal integration mechanism to decode constraints that static baselines perceive as noise. Consequently, we propose that biological efficiency constraints are not computational obstacles but evolutionary pressures that drive the emergence of robust, generalizable intelligence.
@@ Appendix A Appendix
The selection of the modified Duffing oscillator in this study is grounded in prior investigation of dynamical alignment[chen2025dynamical]: A comparative analysis across six classical chaotic attractors (Lorenz, Rössler, Chua, etc.) established that the computational regime of a neural network is primarily governed by the global evolution of phase space volume rather than specific geometric manifolds. The Duffing oscillator mitigates this by providing a unified system where global dynamics can be tuned continuously, thereby isolating the effect of temporal constraint from geometric variation and serving as a standardized probe. Each input feature $x_{i}$ initializes a three-dimensional dynamical system. The initial state $(x_{0},y_{0},z_{0})$ is set as:
@@
To characterize the physical behavior of this system, we systematically mapped its dynamical properties across the control spectrum. Table 4 presents the relationship between $\delta$ and three key dynamical indicators: the Maximum Lyapunov Exponent ($\lambda_{\max}$), the Lyapunov Sum ($\Sigma\lambda_{i}$)[wolf1985determining], and Active Information Storage (AIS)[lizier2012local]. As evidenced by the data, the parameter $\delta$ provides monotonic control over the phase space contraction rate ($\Sigma\lambda_{i}$), spanning from the expansive regime ($\Sigma\lambda_{i}>0$) to the strongly dissipative regime ($\Sigma\lambda_{i}\ll 0$).
Table 4: Dynamical property reference of the mixed oscillator system[chen2025dynamical]. The parameter $\delta$ regulates global phase space dynamics. AIS (Active Information Storage) values in italics highlight the information bottleneck in the transition regime ($\delta\approx 2.0$).

### v2

@@ 2.4 Experiment 3: Zero-Shot Transfer to Unseen Physical Regimes
The preceding experiments established that dissipative constraints embedded in input encoding induce generalization at representational (Exp 1) and structural (Exp 2) levels. This naturally motivates our final test: if low-level invariance and structural emergence can be induced by temporal constraints, does this translate into high-level behavioral robustness in sequential decision-making tasks? To address this, we adopted a zero-shot generalization paradigm in reinforcement learning (RL) tasks [44, 31]: agents trained on a standard “Easy” environment were evaluated on progressively difficult physical regimes without retraining. Unlike approaches relying on extrinsic domain randomization [54, 46], we posit that optimal dynamical constraints confer intrinsic robustness regardless of implementation level. We therefore designed a two-pathway validation:
  • Encoding-Level. To maintain consistency with our previous findings, we first applied the same encoding-level constraints to the CartPole task [55] within a zero-shot paradigm. Dynamical-encoded inputs were processed by SNNs or MLPs to directly isolate the role of temporal integration.
@@
  • Architecture-Level. While dynamical encoding probes how constraints influence learning, it relies on synthetically expanding input vectors into trajectories. However, RL environments provide inherently sequential signals. This enables a direct test of whether the constraint principle generalizes beyond artificial injection. We therefore varied the SNN’s membrane leak parameter $\beta$, which governs intrinsic temporal integration. This parallels our encoding-level constraints: just as $\delta$ controls phase space contraction of input trajectories, $\beta$ controls information dissipation within the network’s hidden states. We conducted systematic $\beta$-sweeps across CartPole (4D state, REINFORCE) and LunarLander [55] (8D state, PPO [51]), comparing Leaky SNNs against MLP and LSTM baselines on raw state histories without external encoding (see Appendix A.3.2 for details).
@@
Both validation approaches reveal consistent patterns (Tables 23): At the encoding level, SNN-Transition ($\delta=2.0$) achieved the lowest generalization gap (38.6 under fixed-budget; 42.4 with sufficient-training), significantly outperforming Expansive (87.6), Dissipative (53.1), and MLP Baseline (112.9). At the architecture level, intermediate dissipation ($\beta=0.5$) consistently achieved optimal transfer: 90% improvement over MLP on CartPole, and 46–50% improvement over MLP and LSTM on LunarLander. Notably, SNNs exhibit non-monotonic performance across the $\beta$ spectrum: over-constrained regimes ($\beta\leq 0.3$) showed training instability, while unconstrained SNNs ($\beta=1.0$) degraded generalization on complex tasks despite successful training. This cross-implementation convergence confirms that the optimal regime reflects a characteristic timescale balancing signal preservation against noise suppression, rather than implementation details.
@@
Table 2: Encoding-level generalization comparison (CartPole). Dynamical encoding transforms states into temporal trajectories with varying constraint strength ($\delta$). SNN-Transition ($\delta=2.0$) achieves the lowest gap. Bold: best generalization (lowest gap).
@@
Table 3: Architecture-level generalization comparison. Varying membrane leak $\beta$ reveals a non-monotonic optimum. Note that over-constrained regimes ($\beta\leq 0.3$) fail to converge reliably. Bold: best generalization (lowest gap).
@@
Furthermore, the non-monotonic performance profile observed in Table 3 helps disentangle a critical confound: does generalization originate from the temporal constraint itself, or merely from the “memory” capacity inherent in SNN membrane dynamics [19]? If memory alone drove generalization, performance should increase monotonically with information retention (high $\beta$). Instead, the observed degradation (or failure) at $\beta\approx 1.0$ indicates that intermediate dissipation implements a necessary temporal information bottleneck, filtering transient noise while preserving invariant structure. This interpretation is further reinforced by the evidence from additional validation with recurrent variants (RLeaky SNNs, see Appendix 16): despite possessing additional memory pathways, these networks did not exhibit improved generalization. On the contrary, they displayed a significantly narrower stability window, failing to converge in unconstrained regimes ($\beta\approx 1.0$) where feedforward Leaky SNNs still functioned. This fragility suggests that excess memory without sufficient constraint destabilizes learning, revealing an embodied duality of the dynamical constraint, as it simultaneously governs both information retention (memory) and phase space contraction (forgetting).
In conclusion, these multi-level validations provide converging evidence for our central thesis: structured dissipation functions not as a limitation but as a temporal inductive bias, compelling networks to extract robust, invariant representations. The consistency of these results across our dual-pathway evaluation confirms that the generalization benefit is an intrinsic property of the constraint principle itself rather than the specificities of implementation.
@@ 3 Discussion
Related Work Recent literature increasingly frames physical constraints not as computational limitations, but as essential inductive biases for intelligence. These constraints can be implemented at different system levels: through architectural design [24], learnable internal dynamics [63], or temporal hierarchies [14, 41]. Our work advances this perspective by identifying temporal dissipation as a unique class of constraint, effective whether imposed via external input dynamics or internal state specification.
As demonstrated in Experiment 2, dissipative dynamics achieve spectral compression equivalent to Slow Feature Analysis [59], but as a physical consequence of phase space contraction rather than an explicit optimization objective. Critically, unlike spatial regularizers where dimensions are interchangeable, temporal constraints exploit an intrinsic asymmetry: the arrow of time transforms irreversible dissipation into a computational resource for extracting invariants. This positions temporal dissipation as a regularization mechanism parallel to spatial constraints: both impose structure that promotes generalization, but through distinct inductive biases. Our results indicate that temporal constraints alone can induce invariant representations required for out-of-distribution generalization [16, 21].
@@
From Generalization to Specialization Our experiments reveal another phenomenon: the transition regime exhibits slower convergence compared to expansive or strongly dissipative dynamics, yet achieves superior generalization. To rigorously ground this observation, we formalize this trade-off phenomenon through a complementary PAC-Bayes analysis [40] (Appendix A.4): by bounding the generalization error as a function of the posterior’s alignment with the prior, this analysis reveals that the transition regime achieves the optimal balance between empirical fit (training error) and structural stability (Kullback–Leibler divergence) while exhibiting the most stable gradient norms (Appendix Table 1819). This provides a principled explanation for the observed cost of abstraction, i.e., the increased training effort: as evidenced by the delayed convergence in our behavioral tasks (Exp 3, Appendix Figure 7), extracting robust invariants is inherently more computationally demanding than the rapid memorization of transient patterns. In other words, the very regime that appears suboptimal from an efficiency-centric view is, in fact, optimal for learning stable, transferable solutions.
@@
This insight not only aligns with the "Edge of Chaos" theory [35, 9] but also explains its computational capacity by explicitly identifying state temporal dissipative constraints as an active regularization mechanism. Taken together, we envision a different landscape: rather than seeking a static optimum for all computation, the neural computation landscape should be viewed as a continuous dynamical spectrum, where the transition regime acts as an adaptability reservoir, maintaining the scale-invariant complexity required to extract robust invariants. Adjacent regimes serve distinct downstream objectives: shifting towards expansive dynamics ($\Sigma\lambda_{i}>0$) maximizes discriminability for inference, or collapsing towards dissipative dynamics ($\Sigma\lambda_{i}\ll 0$) ensures stability and energy minimization. Therefore, this suggests that neural systems are better formalized as adaptive systems rather than static entities. The “optimal” configuration is not a fixed point but a regime to be traversed as task demands change, and neuronal properties become mutable states that are actively evolve to adapt to external environmental impacts and internal budgets.
@@
For machine learning, this necessitates a paradigm shift from architectural search to system dynamical regime navigation: robust generalization emerges from systems that modulate their temporal dynamics in response to task demands between retaining specific data fidelity and extracting generalizable understanding. The transition regime plays the role of a generalization optimum at a certain stage, which emerges from aligning the integration timescale to task-relevant invariants. This perspective offers a potential explanation for the apparent paradox that biological neural systems, despite vast memory capacity or energy efficiency, exhibit robust generalization: sparse coding and metabolic constraints may represent evolutionary solutions that naturally situate neural dynamics within the transition regime.
Thermodynamic Perspective Why should dissipation, typically viewed as loss, confer computational principle? Our findings invite interpretation through non-equilibrium thermodynamics. From the perspective of the Second Law, an unconstrained expansive system naturally drifts toward maximal entropy, which is reflected in our experiments as isotropic noise and failure to generalize. The emergence of structured representations under dissipative constraints can be viewed as the formation of a dissipative structure [47]: by continuously contracting phase space, the network actively resists entropic drift to maintain low-entropy invariant representations.
This perspective provides an explanation for why temporal irreversibility is computationally vital. Unlike reversible Hamiltonian dynamics where information is conserved, intelligence systems require the arrow of time: it is the irreversible loss of transient information (forgetting) that drives the consolidation of stable structure (memory). Dissipative constraints enable networks to compress temporal variability into stable, generalizable representations, much as living systems maintain order through continuous entropy export [50]. From this view, the metabolic constraints observed in biological neural systems are not mere resource limitations, but the thermodynamic conditions under which robust generalization becomes possible.
@@
Limitations and Future Directions While this work demonstrates the efficacy of dynamical constraints, our implementation primarily employs a fixed oscillator system as a controlled probe for mechanistic discovery. This represents only a static instantiation; performance reported herein likely represents a lower bound on the paradigm’s potential. Similarly, although our architecture-level experiments ($\beta$-sweeps) demonstrate that the constraint principle generalizes beyond dynamical encoding, both implementations employ only fixed parameters for the network. Meanwhile, biological systems exhibit heterogeneous time constants across neural populations; exploring learned or adaptive $\beta$ distributions represents a natural extension. Importantly, the transition regime is defined not by a specific parameter values but a spectral signature: high entropy (preserved information complexity) combined with low dominant frequency (alignment with neural networks’ spectral bias). Our scale-space analysis (Fig. 5B) confirms this signature is scale-invariant, suggesting an intrinsic property of the dynamical regime rather than a tuning artifact. Nevertheless, verifying this regime-generality across alternative dynamical systems remains an important direction for future work.
@@ 4 Conclusion
This work identifies dissipative temporal dynamics as a distinct class of inductive bias, effective whether imposed through external or internal states. The critical transition regime does not merely filter noise; it reshapes the learning problem itself, compelling networks to extract invariants that static or unconstrained systems cannot access. These findings challenge a prevailing assumption in machine learning that removing constraints unlocks capability. Our results suggest the opposite relationship: the capacity for generalization may not only exist in architectural complexity or parameter scale, but in the dynamical regime through which information flows. This reframes the design objective from expanding model capacity to structuring the temporal dynamics through which abstraction learning occurs or where transient variation naturally separates from persistent structure. From this perspective, the metabolic pressures that confine biological neural dynamics to dissipative regimes may thus reflect not evolutionary compromise, but a principled solution to the generalization problem.
@@ Appendix A Appendix
The central methodological challenge in studying temporal constraints as inductive biases is to isolate their effect from confounding factors. Classical chaotic systems couple their dynamical properties with specific geometric structures (e.g., the Lorenz butterfly, Rössler spiral, or Chua double-scroll), making it difficult to attribute computational effects to dynamics versus geometry.
We selected a modified duffing oscillator [32] for its minimal geometric complexity. As a damped nonlinear oscillator, the Duffing system produces smooth, non-fractal phase trajectories. This geometric simplicity ensures that observed effects on generalization can be attributed to the global contraction rate ($\Sigma\lambda_{i}$) rather than attractor-specific topology. Concretely, this system offers three methodological advantages: (1) a single parameter $\delta$ permits continuous control of global phase space contraction, spanning from expansive ($\Sigma\lambda_{i}>0$) to strongly dissipative ($\Sigma\lambda_{i}\ll 0$) regimes; (2) this parametric control decouples contraction rate from topological structure; and (3) its well-characterized bifurcation behavior [32] enables principled interpretation across the dynamical spectrum. Prior comparative work across classical chaotic attractors further supports this choice, having established that computational outcomes depend primarily on global Lyapunov structure rather than specific geometric manifolds [60, 13]. This makes the Duffing system an ideal standardized probe for our investigation. Each input feature $x_{i}$ initializes a three-dimensional dynamical system. The initial state $(x_{0},y_{0},z_{0})$ is set as:
@@
To characterize the physical behavior of this system, we systematically mapped its dynamical properties across the control spectrum. Table 4 presents the relationship between $\delta$ and three key dynamical indicators: we employed the QR decomposition method [60] to calculate Lyapunov exponents. At each integration step, we computed the Jacobian matrix $J(\mathbf{x})$, updated an orthonormal basis via $Q_{\text{new}}=e^{J\Delta t}Q$, and performed periodic QR decomposition to accumulate eigenvalue contributions. The Maximum Lyapunov Exponent ($\lambda_{\max}$) and Lyapunov Sum ($\Sigma\lambda_{i}$) were obtained by normalizing accumulated values over total integration time. To characterize encoded trajectory quality, we computed Active Information Storage (AIS) [37], which measures the mutual information between current and preceding states: $\text{AIS}=I(X_{t};X_{t-1})$. Probability densities were estimated using equal-width histograms with Laplace smoothing, and the final metric was averaged across the three spatial dimensions. As evidenced by the data, the parameter $\delta$ provides monotonic control over the phase space contraction rate ($\Sigma\lambda_{i}$), spanning from the expansive regime ($\Sigma\lambda_{i}>0$) to the strongly dissipative regime ($\Sigma\lambda_{i}\ll 0$).
Table 4: Dynamical property reference of the mixed oscillator system[13]. The parameter $\delta$ regulates global phase space dynamics. AIS (Active Information Storage) values in italics highlight the information bottleneck in the transition regime ($\delta\approx 2.0$).